### v1

@@
Abstract
In this paper, we quantify the resilience of nonlinear dynamical systems by studying the increased energy used by all inputs of a system that suffers a partial loss of control authority, either through actuator malfunctions or through adversarial attacks. To quantify the maximal increase in energy, we introduce the notion of an energetic resilience metric. Prior work in this particular setting considers only simple linear models and not general nonlinear dynamical systems. We first characterize the mean value of the control signal in both the nominal and malfunctioning systems, which allows us to approximate the energy in the control. We then obtain a worst-case approximation of this energy for the malfunctioning system, over all malfunctioning inputs. Based on this approximation, we derive bounds on the energetic resilience metric when control authority is lost over one actuator. A simulation example on an academic nonlinear system demonstrates that the metric is useful in quantifying the resilience of the system without significant conservatism, despite the approximations used in obtaining control energies.
1 Introduction
@@
Changes to the input structure of a control system, and in particular actuator faults, also account for $44\%$ of all faults in spacecraft attitude control systems [5, 6]. Building redundancy in actuators [7, 8] or using control reconfiguration schemes [8, 9, 6] are common strategies to address this issue, but can lead to significant costs in control system design and operation. Notably, these faults can prevent a control system from achieving a performance objective such as reachability, safety or optimality with respect to a performance criterion. Moreover, mitigating such issues with classical fault-tolerant, robust or adaptive control strategies is not guaranteed to succeed. Fault-tolerant control only addresses actuator failures limited to reduced effectiveness while still maintaining controllability [10, 11], or actuators “locked” into producing constant inputs [12]. Considering uncontrolled inputs as external disturbances within the paradigm of robust control theory may fail as shown in [13], as robust control requires disturbances to have much smaller magnitudes than the controlled inputs [14], which is not guaranteed. Fast changes in uncontrolled inputs within a short period of time also lead to the failure of adaptive control techniques [14, 15].
@@
Motivated by the difficulties in addressing such failure modes, in this paper we consider nonlinear dynamical systems that suffer from a loss of control authority over a subset of their actuators. Such a failure separates the inputs into controlled and uncontrolled sets. Uncontrolled inputs can potentially be chosen by an adversary and take on any values in their admissible set. However, these values are measurable, and controlled inputs, which are under the authority of the system, can use these measurements to compensate for any undesirable effects from the uncontrolled inputs. In a general nonlinear system, addressing the problem of designing controlled inputs in this setting is difficult. Instead, our objective is to quantify the additional energy in the control inputs used to compensate for the effect of uncontrolled inputs, compared to the nominal case where control authority is not lost. We propose the metric of energetic resilience to quantify this additional energy.
@@
Our line of work is closely related to the work of Bouvier and Ornik [16], who derived conditions for reachability of a target set under any uncontrolled inputs for linear systems, termed resilient reachability. In a subsequent extension [13], a control law that enables a linear system to achieve a target despite an actuator malfunction was designed using the above conditions. In [17], a quantitative resilience metric was introduced by Bouvier et al. for linear driftless systems. This metric was defined as the maximal ratio of minimal reach times to achieve a target state between the nominal and malfunctioning systems. Such a metric quantifies the maximal additional time taken by the malfunctioning system, compared to a nominal system, in achieving its objective. This metric was further used to quantify resilience in generalized integrators in [18], general linear systems in [19, 20] and cyber-physical networks in [21].
@@
In contrast to the above works, the energetic resilience metric we use considers minimal control energies to achieve a target rather than minimal reach times as with quantitative resilience. This metric was first introduced by Padmanabhan et al. [22] to quantify the maximal cost of disturbance for a linear system affected by bounded disturbances. Subsequently, in [22], this metric was used to quantify energetic resilience for the special case of linear driftless systems.
@@
Recent advances have also been made in addressing the issue of resilience for nonlinear systems. De Persis and Tesi [23] and Zhang et al. [24] consider Denial-of-Service attacks on control and cyber-physical systems with nonlinear dynamics. Along similar lines, Pang et al. [25] discuss resilient control for nonlinear cyber-physical systems subjected to network attacks. Zhao et al. [26] and Jia et al. [27] also address the issues of deception and Denial-of-Service attacks over cyber-physical networks. In the problem of resilience for nonlinear systems subjected to external disturbances, Saoud et al. [28] use temporal logic specifications to introduce and quantify a metric for resilience. In a similar setting, Zhou et al. [29] use robust passivity theory to design resilient controllers for nonlinear systems with bounded disturbances. We note that to the best of our knowledge, none of these recent advances address actuator malfunctions or adversaries gaining control over a set of inputs. Our objective is to address resilience in such nonlinear dynamical systems with a set of malfunctioning inputs.
The contributions of this paper are organized as follows. In Section 2, we describe the structure of the nonlinear system we analyze, as well as providing the definition of key quantities derived throughout this paper. In particular, we define a resilience metric which aims to quantify the additional energy required to compensate the effect of malfunctioning inputs. In Section 3, we obtain a characterization of the mean value of the control signal that achieves a specific task in both the nominal and malfunctioning cases. This characterization is used to approximate the minimum control energies to achieve the task in both cases, obtaining an approximate upper bound in the malfunctioning case. Further, we analyze the case when authority is lost over one actuator, obtaining the approximate malfunctioning energy in this particular setting. In Section 4, in this special case, we obtain a bound on the resilience metric which subsequently bounds from above the maximal additional energy required to achieve a ask in the malfunctioning case, compared to the nominal case. We present a simulation example of a nonlinear dynamical system in Section 5. Here, we show that the bound on the resilience metric derived in Section 4 is useful in characterizing the worst-case control energy required under a variety of malfunctioning inputs. In particular, this bound is not significantly conservative despite the approximations used in Section 3 to obtain nominal and malfunctioning energies.
1.1 Notation and Facts
@@
The set $\mathbb{R}^{+}\coloneqq[0,\infty)$ is the set of all non-negative real numbers. For scalars $z\in\mathbb{R}$, define the $\operatorname*{sign}(\cdot)$ function as $\operatorname*{sign}(z)=z/|z|\in\{-1,+1\}$ if $z\neq 0$, with $\operatorname*{sign}(0)=0$. For vectors $z\in\mathbb{R}^{n}$, the $\operatorname*{sign}(\cdot)$ function operates elementwise on $z$. The $p$-norm of a vector $x\in\mathbb{R}^{n}$ is defined as $\|x\|_{p}\coloneqq\left(\sum_{i=1}^{n}|x_{i}|^{p}\right)^{1/p}$, with $\|x\|_{\infty}\coloneqq\max_{i}|x_{i}|$. For a matrix $L\in\mathbb{R}^{p\times q}$ with entries indexed $l_{ij}$, define the induced matrix norms $\|L\|_{1}\coloneqq\max_{j}\sum_{i=1}^{p}|l_{ij}|$ and $\|L\|_{\infty}\coloneqq\max_{i}\sum_{j=1}^{q}|l_{ij}|$. The Moore-Penrose inverse [30], also called the pseudoinverse of $L$, is denoted $L^{\dagger}$. For a piecewise continuous function $u:[0,t_{f}]\to\mathbb{R}^{p}$, its $\mathcal{L}_{2}$ norm is defined as $\|u\|_{\mathcal{L}_{2}}\coloneqq\sqrt{\int_{t=0}^{t_{f}}\|u(t)\|_{2}^{2}~{}\mathrm{d}t}$, and represents the energy in the function over the interval $[0,t_{f}]$. The mean value of such a function is defined as $\overline{u}\coloneqq\frac{1}{t_{f}}\int_{0}^{t_{f}}u(t)\mathrm{d}t$.
@@ 2 Problem Formulation
where the state $x(t)\in\mathbb{R}^{n}$ and the control $u(t)\in\mathbb{R}^{m+p}$ for all $t$, $f:\mathbb{R}^{n}\to\mathbb{R}^{n}$ and $g:\mathbb{R}^{n}\to\mathbb{R}^{n\times(m+p)}$. We assume the functions $f$ and $g$ satisfy the following conditions:
@@
[…]
Condition (2) is similar to a Lipschitz condition on $f$ in the $\infty$-norm. In particular, the Lipschitz constant of $f$ in the $\infty$-norm, if it exists, satisfies (2) for $D_{f}$. Condition (3) bounds $g$ everywhere in $x$. Further, the control $u$ lies in the admissible set $\operatorname*{\mathcal{U}}$, where
$\operatorname*{\mathcal{U}}\coloneqq\left\{u:\mathbb{R}^{+}\to\mathbb{R}^{m+p}:u\text{ is piecewise continuous in $t$, }\|u(t)\|_{\infty}\leq 1\text{ for all $t$}\right\}$ (4)
@@
where the subscripts $c$ and $uc$ denote controlled and uncontrolled respectively. Here, the functions $g_{c}:\mathbb{R}^{n}\to\mathbb{R}^{n\times m}$ and $g_{uc}:\mathbb{R}^{n}\to\mathbb{R}^{n\times p}$ are such that $g(\cdot)=\left[g_{c}(\cdot),~{}g_{uc}(\cdot)\right]$ satisfies (3). Based on the set of admissible controls $u\in\operatorname*{\mathcal{U}}$ with $u(t)=\left[u_{c}^{\operatorname*{\mathsf{T}}}(t),~{}u_{uc}^{\operatorname*{\mathsf{T}}}(t)\right]^{\operatorname*{\mathsf{T}}}$, the controlled and uncontrolled inputs satisfy $u_{c}\in\operatorname*{\mathcal{U}_{c}}$ and $u_{uc}\in\operatorname*{\mathcal{U}_{uc}}$, where
$\operatorname*{\mathcal{U}_{c}}$$\coloneqq\left\{u_{c}:\mathbb{R}^{+}\to\mathbb{R}^{m}:u_{c}\text{is piecewise continuous in $t$, }\|u_{c}(t)\|_{\infty}\leq 1~{}\text{for all $t$}\right\},$ (6)
$\operatorname*{\mathcal{U}_{uc}}$$\coloneqq\left\{u_{uc}:\mathbb{R}^{+}\to\mathbb{R}^{p}:u_{uc}\text{ is piecewise continuous in $t$, }\|u_{uc}(t)\|_{\infty}\leq 1~{}\text{for all $t$}\right\}.$ (7)
In this setting, we consider the task of finite-time reachability, i.e., achieving a specified target state $x(t_{f})=x_{tg}\in\mathbb{R}^{n}$ from an initial state $x(0)=x_{0}\in\mathbb{R}^{n}$ at a specified final time $t_{f}$. Under the nominal dynamics (1), the system uses the input $u\in\operatorname*{\mathcal{U}}$, while under the malfunctioning dynamics (5), the system must use only the controlled input $u_{c}\in\operatorname*{\mathcal{U}_{c}}$ to achieve this task. We note that in the latter case, the uncontrolled input $u_{uc}$ is arbitrarily chosen — potentially by an adversary — in the space $\operatorname*{\mathcal{U}_{uc}}$. However, we assume this choice of $u_{uc}$ is observable, potentially a priori, and hence its effect can be compensated for in the design of $u_{c}$. Achieving this task despite the effect of $u_{uc}$ corresponds to the notion of resilient reachability, formally defined below.
@@
A target $x_{tg}\in\mathbb{R}^{n}$ is resiliently reachable at a given final time $t_{f}$ from $x_{0}\in\mathbb{R}^{n}$ if for all uncontrolled inputs $u_{uc}\in\operatorname*{\mathcal{U}_{uc}}$ there exists a controlled input $u_{c}\in\operatorname*{\mathcal{U}_{c}}$, possibly dependent on $u_{uc}$, such that $x(t_{f})=x_{tg}$.
@@
Characterize the control input $u\in\operatorname*{\mathcal{U}}$ in (1) (resp. $u_{c}\in\operatorname*{\mathcal{U}_{c}}$ in (5)) that ensures a target $x_{tg}\in\mathbb{R}^{n}$ is reachable (resp. resiliently reachable).
In Section 3, we use the solution of the nominal (1) and malfunctioning (5) systems to obtain a condition on the controls $u\in\operatorname*{\mathcal{U}}$ and $u_{c}\in\operatorname*{\mathcal{U}_{c}}$ that achieve finite-time reachability. Further, in both the nominal and malfunctioning cases, we are interested in control signals that use minimum energy to achieve $x(t_{f})=x_{tg}$. In such a context, we make the following definitions, also used in [34].
@@
the dependence of $u_{c}$ on $u_{uc}$ is explicitly denoted with a slight abuse of notation.
In (9), the controlled input $u_{c}$ may depend not only on past and current values of $u_{uc}$, but potentially also future values of this uncontrolled input. The dependence on future values considers any prior knowledge we may have about uncontrolled inputs, such as the corresponding actuators locking in place or acting with reduced effectiveness [11, 12]. The controlled input $u_{c}$ seeks to compensate for the effect of $u_{uc}$ and ensure that $x_{tg}$ is reachable from $x_{0}$ in time $t_{f}$, while also using minimal energy (9). On the other hand, the uncontrolled input $u_{uc}$ also contributes to the energy used by the actuators of the system. In a system (5) that has lost authority over some of its actuators, it is important to characterize the maximal additional energy that may be used by all actuators. In practical cases, this maximal additional energy is directly related to consumption of resources such as fuel in vehicular systems. Quantifying such an increase in resource consumption is essential to understand the maximum capacity of resources required. In an adversarial setting, the uncontrolled input $u_{uc}$ may be chosen to maximize the total energy used by the system. We thus use the following definitions, first introduced in [34], that consider the total energy used by all actuators of the system (5).
@@
Finally, note that the worst-case total energy (11) can be considerably larger than the nominal energy (8) due to the action of the uncontrolled input $u_{uc}$. We thus arrive at our second objective.
@@
To quantify such an increase in energy, we define the following energetic resilience metric, similar to that defined in [22, 34].

### v3

@@
Abstract.
In this paper, we quantify the resilience of nonlinear dynamical systems by studying the increased energy used by all inputs of a system that suffers a partial loss of control authority, either through actuator malfunctions or through adversarial attacks. To quantify the maximal increase in energy, we introduce the notion of an energetic resilience metric. Prior work in this particular setting does not consider general nonlinear dynamical systems. In developing this framework, we first consider the special case of linear driftless systems and recall the energies in the control signal in the nominal and malfunctioning systems. Using these energies, we derive a bound on the energetic resilience metric. For general nonlinear systems, we first obtain a condition on the mean value of the control signal in both the nominal and malfunctioning systems, which allows us to approximate the energy in the control. We then obtain a worst-case approximation of this energy for the malfunctioning system, over all malfunctioning inputs. Assuming this approximation is exact, we derive bounds on the energetic resilience metric when control authority is lost over one actuator. A set of simulation examples demonstrate that the metric is useful in quantifying the resilience of the system without significant conservatism, despite the approximations used in obtaining control energies for nonlinear systems.
This work was supported by Air Force Office of Scientific Research grant FA9550-23-1-0131 and NASA University Leadership Initiative grant 80NSSC22M0070. (Corresponding Author: Ram Padmanabhan. Email: ramp3@illinois.edu.)
1. Introduction
@@
Changes to the input structure of a control system, and in particular actuator faults, also account for $44\%$ of all faults in spacecraft attitude control systems [5, 6]. Building redundancy in actuators [7, 8] or using control reconfiguration schemes [8, 9, 6] are common strategies to address this issue, but can lead to significant costs in control system design and operation. In particular, redundant actuators may suffer from failure of individual components [10] or unintended effects such as counteracting other actuators [11]. Notably, these faults can prevent a control system from achieving a performance objective such as reachability, safety or optimality with respect to a performance criterion. Moreover, mitigating such issues with classical fault-tolerant, robust or adaptive control strategies is not guaranteed to succeed. Fault-tolerant control only addresses actuator failures limited to reduced effectiveness while still maintaining controllability [12, 13], or actuators “locked” into producing constant inputs [14]. Considering uncontrolled inputs as external disturbances within the paradigm of robust control theory may fail as shown in [15], as robust control requires disturbances to have much smaller magnitudes than the controlled inputs [16], which is not guaranteed. Fast changes in uncontrolled inputs within a short period of time also lead to the failure of adaptive control techniques [16, 17].
@@
Motivated by the difficulties in addressing such failure modes, in this paper we consider nonlinear dynamical systems that suffer from a loss of control authority over a subset of their actuators. Such a failure separates the inputs into controlled and uncontrolled sets. Uncontrolled inputs can potentially be chosen by an adversary and take on any values in their admissible set. However, these values are measurable, and controlled inputs, which are under the authority of the system, can use these measurements to compensate for any undesirable effects from the uncontrolled inputs. In a general nonlinear system, addressing the problem of designing controlled inputs in this setting is difficult. Rather than designing these controlled inputs, our objective is to quantify the additional energy in the control inputs used to compensate for the effect of uncontrolled inputs, compared to the nominal case where control authority is not lost. We propose the metric of energetic resilience to quantify this additional energy. We remark that the maximal additional energy is directly related to consumption of resources such as fuel in practical vehicular systems that may lose authority over some inputs. In such settings, fuel economy is an important consideration [18, 19]. Thus, such a quantification of the increased resource consumption can enable an understanding of the maximal capacity required.
@@
Our line of work contrasts with earlier approaches to quantify resilience in systems and networks. Bouvier and Ornik [20] derived conditions for reachability of a target set under any uncontrolled inputs for linear systems, termed resilient reachability. In a subsequent extension [15], a control law that enables a linear system to achieve a target despite an actuator malfunction was designed using the above conditions. In [21], a quantitative resilience metric was introduced by Bouvier et al. for linear driftless systems. This metric was defined as the maximal ratio of minimal reach times to achieve a target state between the nominal and malfunctioning systems. Such a metric quantifies the maximal additional time taken by the malfunctioning system, compared to a nominal system, in achieving its objective. This metric was further used to quantify resilience in generalized integrators in [22], general linear systems in [23, 24] and cyber-physical networks in [25].
@@
In contrast to the above works, the energetic resilience metric we use considers minimal control energies to achieve a target rather than minimal reach times as with quantitative resilience. This metric was first introduced by Padmanabhan et al. [26] to quantify the maximal cost of disturbance for a linear system affected by bounded disturbances. We note that a metric for energetic resilience was proposed in [27] for linear driftless systems. However, that metric is not meaningful for general nonlinear systems, motivating the approach in this paper. We also remark that none of the above works consider a loss of control authority for general nonlinear dynamical systems, and this paper presents our first steps in answering this problem.
@@
Recent advances have been made in addressing the general challenge of resilience for nonlinear systems. De Persis and Tesi [28] and Zhang et al. [29] consider Denial-of-Service attacks on control and cyber-physical systems with nonlinear dynamics. Along similar lines, Pang et al. [30] discuss resilient control for nonlinear cyber-physical systems subjected to network attacks. Zhao et al. [31] and Jia et al. [32] also address the issues of deception and Denial-of-Service attacks over cyber-physical networks. In the problem of resilience for nonlinear systems subjected to external disturbances, Saoud et al. [33] use temporal logic specifications to introduce and quantify a metric for resilience. In a similar setting, Zhou et al. [34] use robust passivity theory to design resilient controllers for nonlinear systems with bounded disturbances. However, to the best of our knowledge, none of these recent advances address actuator malfunctions or adversaries gaining control over a set of inputs. Our objective is to address resilience in such nonlinear dynamical systems with a set of malfunctioning inputs.
The contributions of this paper are organized as follows.
  • In Section 2, we describe the structure of the nonlinear system we analyze, as well as providing the definition of key quantities derived throughout this paper. In particular, we define a resilience metric which aims to quantify the additional energy required to compensate the effect of malfunctioning inputs.
  • In Section 3, we consider the special case of linear driftless systems and recall some relevant results from [27]. Using these results, we briefly discuss a problem of designing the optimal final time to minimize the energy required to achieve a task under a malfunction.
  • Further, in Section 3 we consider the case when control authority is lost over one actuator, deriving the maximal total energy used by the linear driftless system in the malfunctioning case. This is used to derive a bound on the resilience metric.
  • From Section 4, we address general nonlinear dynamical systems. We first obtain a condition on the mean value of the control signal that achieves a specific task in both the nominal and malfunctioning cases. This condition is used to approximate the minimum control energies to achieve the task in both cases, obtaining an approximate upper bound in the malfunctioning case.
  • In Section 4, we also analyze the special case when authority is lost over one actuator, obtaining the approximate malfunctioning energy in this particular setting.
  • In Section 5, under this special case, we obtain a bound on the resilience metric which subsequently bounds from above the maximal additional energy required to achieve a task in the malfunctioning case, compared to the nominal case.
  • We present a set of simulation examples on linear driftless, linear and nonlinear dynamical systems in Section 6. Here, we show that the bounds on the resilience metric derived in Sections 3 and 5 are useful in characterizing the worst-case control energy required under a variety of malfunctioning inputs. In particular, this bound is not significantly conservative for nonlinear systems despite the approximations used in Section 4 to obtain nominal and malfunctioning energies.
1.1. Notation and Facts
@@
The set $\mdmathbb{R}^{+}\coloneqq[0,\infty)$ is the set of all non-negative real numbers. For scalars $z\in\mdmathbb{R}$, define the $\operatorname*{sign}(\cdot)$ function as $\operatorname*{sign}(z)=z/|z|\in\{-1,+1\}$ if $z\neq 0$, with $\operatorname*{sign}(0)=0$. For vectors $z\in\mdmathbb{R}^{n}$, the $\operatorname*{sign}(\cdot)$ function operates elementwise on $z$. The $p$-norm of a vector $x\in\mdmathbb{R}^{n}$ is defined as $\|x\|_{p}\coloneqq\left(\sum_{i=1}^{n}|x_{i}|^{p}\right)^{1/p}$, with $\|x\|_{\infty}\coloneqq\max_{i}|x_{i}|$. For a matrix $L\in\mdmathbb{R}^{p\times q}$ with entries indexed $l_{ij}$, define the induced matrix norms $\|L\|_{1}\coloneqq\max_{j}\sum_{i=1}^{p}|l_{ij}|$ and $\|L\|_{\infty}\coloneqq\max_{i}\sum_{j=1}^{q}|l_{ij}|$. The Moore-Penrose inverse [35], also called the pseudoinverse of $L$, is denoted $L^{\dagger}$. For a piecewise continuous function $u:[0,t_{f}]\to\mdmathbb{R}^{p}$, its $\mathcal{L}_{2}$ norm is defined as $\|u\|_{\mathcal{L}_{2}}\coloneqq\sqrt{\int_{t=0}^{t_{f}}\|u(t)\|_{2}^{2}~\mathrm{d}t}$, and represents the energy in the function over the interval $0,t_{f}]$. The mean value of such a function is defined as $\overline{u}\coloneqq\frac{1}{t_{f}}\int_{0}^{t_{f}}u(t)\mathrm{d}t$.
@@ 2. Problem Formulation
where the state $x(t)\in\mdmathbb{R}^{n}$ and the control $u(t)\in\mdmathbb{R}^{m+p}$ for all $t$, $f:\mdmathbb{R}^{n}\to\mdmathbb{R}^{n}$ and $g:\mdmathbb{R}^{n}\to\mdmathbb{R}^{n\times(m+p)}$. We assume the functions $f$ and $g$ satisfy the following conditions:
$f(x_{1})=f(x_{2})$ $+d_{f}(x_{1},x_{2})$ $\text{ where }\|d_{f}(x_{1},x_{2})\|_{\infty}\leq D_{f}\|x_{1}-x_{2}\|_{\infty},$ (2)
$\text{ and }g(x_{1})=g(x_{2})$ $+d_{g}(x_{1},x_{2})$ $\text{ where }\|d_{g}(x_{1},x_{2})\|_{\infty}\leq D_{g}\|x_{1}-x_{2}\|_{\infty}.$ (3)
In other words, $f$ and $g$ are $D_{f}$ and $D_{g}$-Lipschitz continuous in the $\infty$-norm respectively. Further, the control $u$ lies in the admissible set $\operatorname*{\mathcal{U}}$, where
@@
where the subscripts $c$ and $uc$ denote controlled and uncontrolled respectively. Here, the functions $g_{c}:\mdmathbb{R}^{n}\to\mdmathbb{R}^{n\times m}$ and $g_{uc}:\mdmathbb{R}^{n}\to\mdmathbb{R}^{n\times p}$ are such that $g(\cdot)=\left[g_{c}(\cdot),~g_{uc}(\cdot)\right]$ satisfies (3). Based on the set of admissible controls $u\in\operatorname*{\mathcal{U}}$ with $u(t)=\left[u_{c}^{\operatorname*{\mathsf{T}}}(t),~u_{uc}^{\operatorname*{\mathsf{T}}}(t)\right]^{\operatorname*{\mathsf{T}}}$, the controlled and uncontrolled inputs satisfy $u_{c}\in\operatorname*{\mathcal{U}_{c}}$ and $u_{uc}\in\operatorname*{\mathcal{U}_{uc}}$, where
$\operatorname*{\mathcal{U}_{c}}$ $\coloneqq\left\{u_{c}:\mdmathbb{R}^{+}\to\mdmathbb{R}^{m}:u_{c}\text{ is piecewise continuous in $t$, }\|u_{c}(t)\|_{\infty}\leq 1~\text{for all $t$}\right\},$ (6)
$\operatorname*{\mathcal{U}_{uc}}$ $\coloneqq\left\{u_{uc}:\mdmathbb{R}^{+}\to\mdmathbb{R}^{p}:u_{uc}\text{ is piecewise continuous in $t$, }\|u_{uc}(t)\|_{\infty}\leq 1~\text{for all $t$}\right\}.$ (7)
In this setting, we consider the task of fixed-time reachability, i.e., achieving a specified target state $x(t_{f})=x_{tg}\in\mdmathbb{R}^{n}$ from an initial state $x(0)=x_{0}\in\mdmathbb{R}^{n}$ at a specified final time $t_{f}$. Under the nominal dynamics (1), the system uses the input $u\in\operatorname*{\mathcal{U}}$, while under the malfunctioning dynamics (5), the system must use only the controlled input $u_{c}\in\operatorname*{\mathcal{U}_{c}}$ to achieve this task. We note that in the latter case, the uncontrolled input $u_{uc}$ is arbitrarily chosen — potentially by an adversary — in the space $\operatorname*{\mathcal{U}_{uc}}$. However, we assume this choice of $u_{uc}$ is observable, potentially a priori, and hence its effect can be compensated for in the design of $u_{c}$. Achieving this task despite the effect of $u_{uc}$ corresponds to the notion of resilient reachability, formally defined below.
@@
A target $x_{tg}\in\mdmathbb{R}^{n}$ is resiliently reachable at a given final time $t_{f}$ from $x_{0}\in\mdmathbb{R}^{n}$ if for all uncontrolled inputs $u_{uc}\in\operatorname*{\mathcal{U}_{uc}}$ there exists a controlled input $u_{c}\in\operatorname*{\mathcal{U}_{c}}$, possibly dependent on $u_{uc}$, such that $x(t_{f})=x_{tg}$.
@@
Characterize the control input $u\in\operatorname*{\mathcal{U}}$ in (1) (resp. $u_{c}\in\operatorname*{\mathcal{U}_{c}}$ in (5)) that ensures a target $x_{tg}\in\mdmathbb{R}^{n}$ is reachable (resp. resiliently reachable).
In Section 4, we use the solution of the nominal (1) and malfunctioning (5) systems to obtain a condition on the controls $u\in\operatorname*{\mathcal{U}}$ and $u_{c}\in\operatorname*{\mathcal{U}_{c}}$ that achieve fixed-time reachability. Further, in both the nominal and malfunctioning cases, we are interested in control signals that use minimum energy to achieve $x(t_{f})=x_{tg}$. In such a context, we make the following definitions, also used in [27].
@@
where the dependence of $u_{c}$ on $u_{uc}$ is explicitly denoted with a slight abuse of notation.
In (9), the controlled input $u_{c}$ may depend not only on past and current values of $u_{uc}$, but potentially also future values of this uncontrolled input. The dependence on future values considers any prior knowledge we may have about uncontrolled inputs, such as the corresponding actuators locking in place or acting with reduced effectiveness [13, 14]. In both (8) and (9), deriving the corresponding optimal control signals is not tractable [39] due to the nonlinear nature of the systems (1) and (5). However, our objective in this paper is not to design these optimal control signals, but rather to quantify how much larger the malfunctioning energy may be, compared to the nominal energy.
The controlled input $u_{c}$ seeks to compensate for the effect of $u_{uc}$ and ensure that $x_{tg}$ is reachable from $x_{0}$ in time $t_{f}$, while also using minimal energy (9). On the other hand, the uncontrolled input $u_{uc}$ also contributes to the energy used by the actuators of the system. In a system (5) that has lost authority over some of its actuators, it is important to characterize the maximal additional energy that may be used by all actuators. In practical cases, this maximal additional energy is directly related to consumption of resources such as fuel in vehicular systems. Quantifying such an increase in resource consumption is essential to understand the maximum capacity of resources required. In an adversarial setting, the uncontrolled input $u_{uc}$ may be chosen to maximize the total energy used by the system. We thus use the following definitions, first introduced in [27], that consider the total energy used by all actuators of the system (5).
@@
Finally, note that the worst-case total energy (11) can be considerably larger than the nominal energy (8) due to the action of the uncontrolled input $u_{uc}$. Quantifying this increased energy is an important problem to understand additional consumption of resources in practical settings. We thus arrive at our second objective, that of quantifying this additional energy for general nonlinear systems.
@@
To quantify such an increase in energy, we define the following energetic resilience metric, similar to that defined in [26].